OpenStudy (anonymous):

Find the derivative of the function. sqrt(x) + 1/4(sin(2x)^2)

OpenStudy (anonymous):

\[\sqrt{x}+1/4\sin(2x)^{2}\]

OpenStudy (anonymous):

1/2x^-1/2 + 1/2sin(2x) * cos(2x) * 2

OpenStudy (anonymous):

\[1/\sqrt{x} + \sin(2x)\cos(2x)\]

OpenStudy (kinggeorge):

\[{d\over dx} \;\;\; \sqrt x+{1\over4}\sin(2x)^2 = {1\over{2\sqrt x}}+ {2\over2}\sin(2x)\cos(2x)\]

OpenStudy (anonymous):

well the book says

OpenStudy (kinggeorge):

Remember you're also multiplying the \[{1\over \sqrt{x}}\]by \(1\over2\)

OpenStudy (anonymous):

\[1/\sqrt{x} + 2xcos(2x)^{2}\]

OpenStudy (anonymous):

i mean yes the 1/2

OpenStudy (anonymous):

is there an identity that makes it 2xcos(2x)^2

OpenStudy (kinggeorge):

Not that I know of. Let me see if I can get wolfram to change it.

OpenStudy (kinggeorge):

Well, when I use wolfram to integrate \({1\over \sqrt{x}}+2x \cos^2 (2x)\) I don't get the original function back, so I'm thinking the book is wrong.

OpenStudy (anonymous):

k i was wondering myself